OpenStudy (animelover98):

When solving a radical equation, Beth and Kelly came to two different conclusions. Beth found a solution, while Kelly's solution did not work in the equation. Create and justify two situations: one situation where Beth is correct and a separate situation where Kelly is correct.

OpenStudy (animelover98):

@SolomonZelman Could you please help me? Thanks!

OpenStudy (kaylak12345):

i mean -18 and -6

OpenStudy (solomonzelman):

you want a radical equation that has a solution for part A, then another equation with no solution for part B. What the equations are is up to you.

OpenStudy (animelover98):

@kaylak12345 Could you please post that in your own question? Thanks

OpenStudy (solomonzelman):

Can I call you lover?

OpenStudy (animelover98):

@SolomonZelman So I need to get an extraneous solution. And sure, I'll respond to anything haha

OpenStudy (solomonzelman):

yes, extraneous solution... a solution that you arrive at at the end of the solving process but doesn't work.

OpenStudy (solomonzelman):

saying doesn't work in the original equation.

OpenStudy (solomonzelman):

This all is for part B. When Kelly is correct.

OpenStudy (solomonzelman):

For part a, you just need a good solvable radical equation.

OpenStudy (solomonzelman):

For extraneous equation, I would put something like \(\Large\color{black}{ \frac{1}{x+2}+\frac{1}{x-2} =\frac{4}{x^2-4} }\)

OpenStudy (animelover98):

Rather..... \[(\sqrt{9})^2=x^2\]

OpenStudy (solomonzelman):

try to solve this one, and tell me what your solution is. Then check your answer and tell me if it worked.

OpenStudy (solomonzelman):

Ohh, radical equation... sorry.

OpenStudy (animelover98):

No! It's alright! So, do I try that equation or the one you used?

OpenStudy (solomonzelman):

No it is not a radical equation... it is irrelevant. although it does have an extraneous solution.

OpenStudy (animelover98):

Alright, Thank you so much. I have all day so please do not feel rushed.

OpenStudy (solomonzelman):

can one solution be extraneous, and another be not extraneous?

OpenStudy (solomonzelman):

I mean 1 extraneous and another that works. can you use that?

OpenStudy (animelover98):

I believe so

OpenStudy (solomonzelman):

\(\Large\color{black}{ -\sqrt[]{x-1}=x-7 }\)

OpenStudy (solomonzelman):

try to solve this one.

OpenStudy (solomonzelman):

(start from raising both sides to the second power. saying- SQUARE BOTH SIDES)

OpenStudy (animelover98):

\[x-1=x^2-49\]?

OpenStudy (solomonzelman):

No exactly.

OpenStudy (solomonzelman):

you got the left side correctly. and you squared the minus as well, so it is positive x-1 as you wrote. Left side is good.\ When you square the right side, you are taking a square (second power) of \(\Large\color{black}{ (x-7) }\) saying that this is what you do, \(\Large\color{black}{ (x-7) ^{\color{red}{2}}}\) .

OpenStudy (solomonzelman):

what you wrote for the right side is missing the middle term.

OpenStudy (animelover98):

Ah \[x^2-14x-49\]

OpenStudy (animelover98):

I'll keep working. Thanks for your help!

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

So you get \(\Large\color{blue}{ x-1=x^2-14x-49 }\)

OpenStudy (animelover98):

Yes, so now I solve this equation?

OpenStudy (solomonzelman):

yes, I would just add like terms...

OpenStudy (solomonzelman):

-x on both sides, +1 on both sides and on....

OpenStudy (animelover98):

-x on both sides?

OpenStudy (animelover98):

I'm very stuck and confused now. I have no idea what I'm doing.

OpenStudy (solomonzelman):

\(\Large\color{blue}{x-1=x^2-14x-49 \color{red}{ } }\) \(\Large\color{blue}{x-1\color{red}{-x}=x^2-14x-49 \color{red}{-x} }\) \(\Large\color{blue}{-1=x^2-13x-49 }\) \(\Large\color{blue}{-1\color{red}{+1}=x^2-13x-49\color{red}{+1} }\)

OpenStudy (solomonzelman):

it should be -15x on the left, because I subtract -14x -x

OpenStudy (animelover98):

Oh good. I got that when I finished solving it. It just looked so weird I thought I was wrong.

OpenStudy (solomonzelman):

\(\Large\color{black}{x^2-15x-48=0 }\)

OpenStudy (solomonzelman):

Okay, and what are you solutions?

OpenStudy (solomonzelman):

I am wrong again.

OpenStudy (solomonzelman):

it was positive 49, not negatgive, so wen I add 1 to both sides, I get, \(\Large\color{black}{x^2-15x+50=0 }\)

OpenStudy (solomonzelman):

I am so confusing.

OpenStudy (solomonzelman):